If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Answer by CharStar(110)

You would use the formula for work which is
1/rate + 1/rate = 1/time
1/4 + 1/6 = 1/t
Add the fraction together. First you must find a common denominator which is 12. 12 goes into 4 - 3 times and 12 goes into 6 2 times.
3/12 + 2/12 = 5/12= 1/t
Cross Multiply 5/12=1/7 (proportion)
5t=12
The answer will be 2 2/5 hours